Calculate the electric potential and magnitude of electric field strength 0.378 m from a small sphere with 6.0E10 excess electrons.




1
Expert's answer
2021-06-01T10:51:07-0400
"V=k\frac{q}{r}=-9\cdot10^{9}\frac{(6\cdot10^{10})1.6\cdot10^{-19}}{0.378}=-230\ V"

"E=k\frac{q}{r^2}=-9\cdot10^{9}\frac{(6\cdot10^{10})1.6\cdot10^{-19}}{0.378^2}=-600\frac{V}{m}"